# A nonparametric approach to practical identifiability of nonlinear mixed effects models

### v1

@@ 2.3 Analysis of population parameter distributions
We begin by evaluating if the $N$ individual estimates, $\{\phi_{i,j}\}_{i=1}^{N}$, are sufficient to distinguish between the underlying population distributions from two distinct fits. The Kolmogorov-Smirnov two-sample test is a nonparametric test to assess if these $N$ samples were taken from different underlying population distributions. Thus, we consider the individual estimates from each possible pair of distinct fits, $\{\phi_{i,j}\}_{i=1}^{N}$ and $\{\phi_{i,k}\}_{i=1}^{N}$, and use the Kolmogorov-Smirnov test to identify significant differences between $\Phi_{i,j}$ and $\Phi_{i,k}$. We use the significance threshold $\alpha=0.05$ to reject the null-hypothesis that the samples are drawn from the same underlying distribution.
@@
An overlapping index of $o_{i,j,k}=1$ indicates that the two probability measures are equal almost-surely. Conversely, $o_{i,j,k}=0$ implies that the two distributions are entirely disjoint. The overlapping index has recently been used to distinguish between responders and non-responders in a virtual clinical trial [48]. Essentially, this nonparametric approach measures the agreement between the two probability distributions and is directly related to the total variation distance between the corresponding measures.
@@ 3.1 The Friberg model of chemotherapy-induced neutropenia is identifiable at the population level
Fit Number $-2\mathcal{LL}$ Fit Number $-2\mathcal{LL}$ Fit Number $-2\mathcal{LL}$
@@
The goodness-of-fits for these 10 best fits identify three distinct local maxima of the log-likelihood, at $-2\mathcal{LL}\approx-1250,-950$ and 790.-790.- 790 . Despite these differences, the model fits the synthetic data quite well for each of the 10 best fits. Now, we considered a large number of fits from throughout the plausible parameter range. Therefore, we conclude that fits 1-3 from Table 3, with the highest log-likelihood values, represent the global maximum of the log-likelihood, while the other two clusters represent local maxima. For each of the 10 best fits, we show the individual fits to the synthetic circulating neutrophil concentrations in Figure 2. For some individuals, the predicted dynamics are identical across all 10 best fits. However, some individual fits illustrate subtle differences in the predicted neutrophil dynamics during chemotherapy. Specifically, for individuals with oscillatory neutrophil dynamics with a relatively large amplitude (ID: 4, 5, 7, 9, 14), there is a discernible difference between the model predictions across the 10 best fits. These differences are reflected in the corresponding log-likelihood value for each of these fits as shown in Table 3.
@@
Having established that the Friberg model is able to recapture the synthetic data, we next turn to the estimated model parameters. In Figure 3A, we show the distribution of the 15 individual estimates for each of the 10 best fits. For the model parameters $N_{0}$ and $k_{tr}$, the distributions of the individual estimates are similar between all 10 best fits. In contrast, for $k_{prol}$ and $\log(EC_{50}),$ there is a clear clustering of the individual distributions of these parameters for fits (1-3), (4-7), and (8-10), which correspond to the three local maxima of the log-likelihood. As we show in Figure 3B, the Kolmogorov-Smirnov test did not identify any significant pairwise differences between the individual estimates for the 10 best fits of the model for $N_{0}$ and $k_{tr}$. However, there are significant differences, measured via the Kolmogorov-Smirnov test, for the parameters $k_{prol}$ and $\log(EC_{50})$ corresponding to the three local maxima. In Figure 3B, we highlight these clusters of individual estimates. Importantly, while we do identify significant differences between the individual estimates for these parameters at each local maxima, the estimates within each cluster do not significantly differ from each other, which suggests that these are three distinct local optima of the log-likelihood.
[…]
@@ 3.2 The standard viral dynamics model is not identifiable at the population level
Fit $-2\mathcal{LL}$ Fit $-2\mathcal{LL}$ Fit $-2\mathcal{LL}$ Fit $-2\mathcal{LL}$ Fit $-2\mathcal{LL}$
@@
Table 4: Summary of the 10 best fits to the synthetic viral load data. The 10 best fits to the synthetic viral load data using Eq. (3) as measured by the log-likelihood. Here, lower values of $-2\mathcal{LL}$ correspond to a better fit.
We note that all 10 best fits to the synthetic viral load data have extremely similar log-likelihood values, which indicates no appreciable differences between the best-fit and the 10th best fit. Indeed, the worst of the 100 distinct parameterizations of this model has a $-2\mathcal{LL}$ value of -189.95, which is comparable to the best fit. Consequently, we conclude that these 10 best fits have converged to the maximum of the log-likelihood. In Figure 4, we show the individual fits to each of the 15 virtual participants from all 10 best fits. As we would expect from the log-likelihood, the model fits the synthetic data well for each of the 10 best fits. Furthermore, the predicted individual dynamics are indistinguishable for all 15 virtual participants. We therefore conclude that the model can recapture the synthetic data.
@@
We next analyze the parameter estimates from the 10 best fits. Here, we do not show the distribution of $V(0),$ as this parameters is directly informed by the synthetic data. In Figure 5, we show the individual estimates and the population distributions for the model parameters $\log(\beta),\ \log(T_{0}),\ \log(p),$ and $\delta$. The distributions of individual estimates for the parameters $\log\beta$ and $\delta$ are indistinguishable across the 10 best-fits. Accordingly, the Kolmogorov-Smirnov two-sample test does not indicate that there is a significant difference between the individual estimates for these two model parameters. Conversely, the distributions of individual parameter estimates for $\log T_{0}$ and $\log p$ are observably distinct in Figure 5A. The Kolmogorov-Smirnov test identifies significant differences between almost all pairs of individual estimates for these two parameters. In Figure 5B, the large number of blue squares correspond to pairs of individual estimates with $p<0.05$. Consequently, our analysis of the individual estimates suggests that $\log T_{0}$ and $\log p$ may not be identifiable at the population level, as these statistically different distributions result in identical viral load dynamics. This conclusion is consistent earlier structural and practical identifiability analysis of the standard viral dynamics model [46, 75] which showed that $p$ cannot be identified from viral load data alone.
[…]
@@
We next consider the population distributions, which reflect the conclusions from the individual estimates. Once again, the probability densities of population distributions of $\log\beta$ and $\delta$ are indistinguishable in Figure 5C. As expected from the probability densities for the population distribution of these parameters shown in Figure 5C, the overlapping indices for the 10 best fits of $\log\beta$ and $\delta$ are all greater than 0.97 and 0.98, respectively (Figure 5D). Thus, the population distributions for these two parameters have considerable overlap with a correspondingly small total variation distance. Conversely, there is noticeable heterogeneity in the mean and shape of the population distributions of $\log T_{0}$ and $\log p$. This heterogeneity is reflected in the overlapping indices of the population distributions of these two parameters, where the minimum overlap between two distributions is roughly $10^{-3}$ in both cases, which indicates a large total variation distance between the population parameter distributions. Thus, our analysis demonstrates that distinct population distributions for $\log T_{0}$ and $\log p$ can result in nearly identical model predictions and fits to data. This analysis indicates that $\log\beta$ and $\delta$ are identifiable at the population level, while $\log T_{0}$ and $\log p$ are unidentifiable.
@@ 4 Discussion
Nonlinear mixed-effects models are extensively used within the pharmacokinetics modelling community [73, 39, 14] and hierarchical models are increasingly common throughout mathematical biology [36, 27, 8]. Although this modelling framework benefits from simultaneously considering data from multiple individual participants when estimating parameters, existing approaches to study the identifiability of model parameters relies on a parametric interpretation of the population parameter distributions. Here, we have presented an alternative approach to studying the practical identifiability of a nonlinear mixed-effects model through a nonparametric approach. Specifically, rather than studying when the parametric representation of two distributions are precisely equal, we focus on determining when two distributions are statistically different.
@@
We illustrated our approach by analysing the identifiability of two important models in pharmacology and viral dynamics. We showed that the Friberg model of chemotherapy-induced neutropenia is identifiable within a hierarchical framework without imposing that $k_{prol}=k_{tr}.$ Thus, our work demonstrates that considering multiple individuals allows for the Friberg model to be parameterized without assuming that $k_{tr}=k_{prol}$ [24, 5]. Conversely, our results indicate that the standard viral dynamics model remains unidentifiable, even when considering multiple individuals. This population level analysis is consistent with earlier work that considered the identifiability of the standard viral dynamics model when fitting viral load data from a single individual [75, 7]. Consequently, the standard viral dynamics model is an example of where considering multiple individuals does not permit the identification of additional model parameters. Further, our analysis of the standard viral dynamics model indicates that a multi-start fitting approach, as suggested by Duchesne et al., [22], may not be sufficient to determine a unique global minimum, as we found over 10 local minima of the log-likelihood that have nearly identical goodness-of-fits.
We have proposed a nonparametric approach to studying practical identifiability within the framework of nonlinear mixed-effects models. While the proposed approach is simple, it is computationally demanding as we utilize a multi-start approach to identify the global optima of the log-likelihood. Further complicating this multi-start approach, it is not a priori apparent how many runs are necessary to find the global optimum. Further, we proposed using the overlapping index and the Kolmogorov-Smirnov test to determine if the resulting population parameter distributions differ. We focused on characterizing the population distribution of individual parameters, rather than multivariate joint distributions that would arise from an underlying covariate structure. Consequently, there are several areas for future work from both theoretical and practical perspectives. For example, this proposed approach to practical identifiability could be implemented within the convergence assessment framework of existing software packages [63, 8] or the profile likelihood approach to parameter identifiability [60].

### v2

@@ 2.3 Analysis of population parameter distributions
We begin by evaluating if the $N$ individual estimates, $\{\phi_{i,j}\}_{i=1}^{N}$, are sufficient to distinguish between the underlying population distributions from two distinct fits. The Kolmogorov-Smirnov two-sample test is a nonparametric test to assess if these $N$ samples were taken from different underlying population distributions. Thus, we consider the individual estimates from each possible pair of distinct fits, $\{\phi_{i,j}\}_{i=1}^{N}$ and $\{\phi_{i,k}\}_{i=1}^{N}$, and use the Kolmogorov-Smirnov test to identify significant differences between $\Phi_{i,j}$ and $\Phi_{i,k}$. We use the significance threshold $\alpha=0.05$ to reject the null-hypothesis that the samples are drawn from the same underlying distribution. However, the Kolomogorov-Smirnov test relies on the ability to well-order the individual estimates $\Phi_{i,j}$. Consequently, when considering correlated parameters, we consider the nonparametric maximum mean discrepancy test [32, 28] to test if two multivariate samples are drawn from different underlying distributions.
@@
An overlapping index of $o_{i,j,k}=1$ indicates that the two probability measures are equal almost-surely. Conversely, $o_{i,j,k}=0$ implies that the two distributions are entirely disjoint. The overlapping index has recently been used to distinguish between responders and non-responders in a virtual clinical trial [57]. Essentially, this nonparametric approach measures the agreement between the two probability distributions and is directly related to the total variation distance between the corresponding measures. The overlapping index immediately extends to the multivariate case by calculating a higher-dimensional integral and facilitates the comparison of multivariate population parameter distributions.
Finally, model parameters that were estimated at the population-level, i.e. without inter-individual variability, correspond to point-estimates. Consequently, studying the identifiability of these parameters without inter-individual variability does not require understanding if two distributions are significantly different, but rather studying if the data is sufficient to constrain the point-estimate. Therefore, the identifiability of these point-estimates can be directly studied using existing techniques focused on individual-level dynamics such as profile likelihood analysis [70].
@@ 3.1 The Friberg model of chemotherapy-induced neutropenia is identifiable at the population level
Fit Number $2\mathcal{LL}$ Fit Number $2\mathcal{LL}$ Fit Number $2\mathcal{LL}$
@@
The goodness-of-fits for these 10 best fits identify three distinct local maxima of the log-likelihood, at $2\mathcal{LL}\approx-1250,-950$ and 790.-790. Despite these differences, the model fits the synthetic data quite well for each of the 10 best fits. Now, we considered a large number of fits from throughout the plausible parameter range. Therefore, we conclude that fits 1-3 from Table 3, with the highest log-likelihood values, represent the global maximum of the log-likelihood, while the other two clusters represent local maxima. For each of the 10 best fits, we show the individual fits to the synthetic circulating neutrophil concentrations in Figure 2. For some individuals, the predicted dynamics are identical across all 10 best fits. However, some individual fits illustrate subtle differences in the predicted neutrophil dynamics during chemotherapy. Specifically, for individuals with oscillatory neutrophil dynamics with a relatively large amplitude (ID: 4, 5, 7, 9, 14), there is a discernible difference between the model predictions across the 10 best fits. These differences are reflected in the corresponding log-likelihood value for each of these fits as shown in Table 3.
@@
Having established that the Friberg model is able to recapture the synthetic data, we next turn to the estimated model parameters. In Figure 3A, we show the distribution of the 15 individual estimates for each of the 10 best fits. For the model parameters $N_{0}$ and $k_{tr}$, the distributions of the individual estimates are similar between all 10 best fits. In contrast, for $k_{prol}$ and $\log(EC_{50}),$ there is a clear clustering of the individual distributions of these parameters for fits (1-3), (4-7), and (8-10), which correspond to the three local maxima of the log-likelihood. As we show in Figure 3B, the Kolmogorov-Smirnov test did not identify any significant pairwise differences between the individual estimates for the 10 best fits of the model for $N_{0}$ and $k_{tr}$. However, there are significant differences, measured via the Kolmogorov-Smirnov test, for the parameters $k_{prol}$ and $\log(EC_{50})$ corresponding to the three local maxima. In Figure 3B, we highlight these clusters of individual estimates. Importantly, while we do identify significant differences between the individual estimates for these parameters at each local maxima, the estimates within each cluster do not significantly differ from each other, which suggests that these are at least three distinct local optima of the log-likelihood; we can identify either 3 local optima by using the clusters shown in Figure 3, or 4 local optima by dividing the second cluster into two subgroups of fits 4 and 6 and fits 5 and 7. Specifically, there are significant pairwise differences between the individual estimates for $\log_{10}(EC_{50})$ these subgroups. However, the population distributions for these parameters display significant overlap, which implies that the 15 individuals considered in this example may not be sufficient to distinguish between the population estimates for this parameter.
Figure 3: Assessing practical identifiability of the Friberg model. Panel A shows the distribution of the 15 individual estimates for the model parameters $N_{0},k_{tr},k_{prol}$ and $\log(EC_{50})$ for each of the 10 best population fits as dots within each violin plot. Panel B shows the $p$-value corresponding to the pairwise Kolmogorov-Smirnov test between all 10 best fits for each of the four model parameters, with the grey and white borders identifying the three distinct clusters. Here, $p$-values below the significance threshold of $0.05$ are plotted as blue squares. Panel C shows the 10 estimated population distributions for each of the four model parameters. Panel D shows the overlapping index between each of the pairwise comparisons of the 10 best fits. Here, for each of the $k=1,2,3,4$ model parameters, pairs of fits $(i,j)$ with an overlapping index $o_{i,j,k}>0.5$ are represented by the shades of violet, while pairs with $o_{i,j,k}<0.5$ are represented by shades of orange and the three clusters are again highlighted by white boxes.
@@
We next compared the estimates for the parameters considered without inter-individual variability, $\gamma$ and $k_{circ}$, against the true values for these parameters for the cluster of best-fits (1-3) . While we did not precisely estimate the known values, this fitting approximately recaptured the underlying truth values with $\gamma=0.15,0.15,$ and $\gamma=0.18$ for fits 1-3, respectively compared against the true value of $\gamma=0.163$. Similarly, the estimates for $k_{circ}=0.87$/day, 0.810.81/day, and $k_{circ}=1.41$/day compared against $k_{circ}=1.15$/day We also compared the population parameter distributions from fits 1-3 against the known distributions for the parameters with inter-individual variability in Figure 8. In general, the population estimates recaptured the underlying known parameter distributions.
@@ 3.2 The standard viral dynamics model is not identifiable at the population level
Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$
@@
Table 4: Summary of the 10 best fits to the synthetic viral load data. The 10 best fits to the synthetic viral load data using Eq. (3) as measured by the log-likelihood. Here, lower values of $2\mathcal{LL}$ correspond to a better fit.
We note that all 10 best fits to the synthetic viral load data have extremely similar log-likelihood values, which indicates no appreciable differences between the best-fit and the 10th best fit. Indeed, the worst of the 100 distinct parameterizations of this model has a $2\mathcal{LL}$ value of -189.95, which is comparable to the best fit. Consequently, we conclude that these 10 best fits have converged to the maximum of the log-likelihood. In Figure 4, we show the individual fits to each of the 15 virtual participants from all 10 best fits. As we would expect from the log-likelihood, the model fits the synthetic data well for each of the 10 best fits. Furthermore, the predicted individual dynamics are indistinguishable for all 15 virtual participants. We therefore conclude that the model can recapture the synthetic data.
@@
We next analyze the parameter estimates from the 10 best fits. Here, we do not show the distribution of $V(0),$ as this parameters is directly informed by the synthetic data. In Figure 5, we show the individual estimates and the population distributions for the model parameters $\log_{10}(\beta),\ \log_{10}(T_{0}),\ \log_{10}(p),$ and $\delta$. The distributions of individual estimates for the parameters $\log_{10}\beta$ and $\delta$ are indistinguishable across the 10 best-fits. Accordingly, the Kolmogorov-Smirnov two-sample test does not indicate that there is a significant difference between the individual estimates for these two model parameters. Conversely, the distributions of individual parameter estimates for $\log_{10}T_{0}$ and $\log_{10}p$ are observably distinct in Figure 5A. The Kolmogorov-Smirnov test identifies significant differences between almost all pairs of individual estimates for these two parameters. In Figure 5B, the large number of blue squares correspond to pairs of individual estimates with $p<0.05$. Consequently, our analysis of the individual estimates suggests that $\log_{10}T_{0}$ and $\log_{10}p$ may not be identifiable at the population level, as these statistically different distributions result in identical viral load dynamics. This conclusion is consistent with earlier structural and practical identifiability analysis of the standard viral dynamics model [55, 87] which showed that $p$ cannot be identified from viral load data alone.
Figure 5: Assessing practical identifiability of the standard viral dynamics model. Panel A shows the distribution of the 15 individual estimates for the model parameters $\log_{10}(\beta),\log_{10}(T_{0}),\log_{10}(p)$ and $\delta$ for each of the 10 best population fits as dots within each violin plot. Panel B shows the $p$-value corresponding to the pairwise Kolmogorov-Smirnov test between all 10 best fits for each of the four model parameters. Here, $p$-values below the significance threshold of $0.05$ are plotted as blue squares. Panel C shows the 10 estimated population distributions for each of the four model parameters. Panel D shows the overlapping index between each of the pairwise comparisons of the 10 best fits. Here, for each of the $k=1,2,3,4$ model parameters, pairs of fits $(i,j)$ with an overlapping index $o_{i,j,k}>0.5$ are represented by the shades of violet, while pairs with $o_{i,j,k}<0.5$ are represented by shades of orange. The 10 best fits are distinguished by their colors for the distributions shown in Panels A and C.
@@
We next consider the population distributions, which reflect the conclusions from the individual estimates. Once again, the probability densities of population distributions of $\log_{10}\beta$ and $\delta$ are indistinguishable in Figure 5C. As expected from the probability densities for the population distribution of these parameters shown in Figure 5C, the overlapping indices for the 10 best fits of $\log_{10}\beta$ and $\delta$ are all greater than 0.97 and 0.98, respectively (Figure 5D). Thus, the population distributions for these two parameters have considerable overlap with a correspondingly small total variation distance. Conversely, there is noticeable heterogeneity in the mean and shape of the population distributions of $\log_{10}T_{0}$ and $\log_{10}p$. This heterogeneity is reflected in the overlapping indices of the population distributions of these two parameters, where the minimum overlap between two distributions is roughly $10^{-3}$ in both cases, which indicates a large total variation distance between the population parameter distributions. Thus, our analysis demonstrates that distinct population distributions for $\log_{10}T_{0}$ and $\log_{10}p$ can result in nearly identical model predictions and fits to data. This analysis indicates that $\log\beta$ and $\delta$ are identifiable at the population level, while $\log_{10}T_{0}$ and $\log_{10}p$ are unidentifiable.
3.2.1 Identifiability of the standard viral dynamics model with a correlation structure
Further analysis of the identifiability results for the standard viral dynamics model suggest a correlation between $\log T_{0}$ and $\log p$. This result is unsurprising from the biological interpretation of these parameters, as $\log T_{0}$ determines the cells available to be infected, while $\log p$ determines the viral production rate. For a given set-point viral load, corresponding to the equilibrium virus concentration observed at later time points, there is a clear relationship between these two mechanisms. For example, fewer infected cells, with a corresponding lower value of $\log T_{0}$, and a higher production rate $\log p$ can produce the same total amount of virus as more infected cells with a lower production rate, as discussed by Wu et al., [87]. Further, Lavielle and Aarons, [45] has shown that some otherwise unidentifiable parameters may be identifiable within a hierarchical framework with an appropriate correlation structure. We therefore allowed for a correlation between these two parameters and once again fit Eq. (3) to the synthetic viral load from $N=15$ individuals. As before, we considered the 10 best fits to this synthetic data, with the corresponding population log-likelihoods given in Table 5. We once again note that these 10 bests fit the synthetic data equally well, as measured by a negligible difference in the population $2\mathcal{LL}$ values, and all 10 fits capture the individual dynamics.
Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$ Fit $2\mathcal{LL}$
Fit 1 -217.8 Fit 2 -217.8 Fit 3 -217.7 Fit 4 -217.7 Fit 5 -217.7
Fit 6 -217.6 Fit 7 -217.6 Fit 8 -217.6 Fit 9 -217.6 Fit 10 -217.6
Table 5: Summary of the 10 best fits to the synthetic viral load data. The 10 best fits to the synthetic viral load data using Eq. (3) and allowing for correlation between $\log T_{0}$ and $\log p$, as measured by the log-likelihood. Here, lower values of $2\mathcal{LL}$ correspond to a better fit.
We analyse the parameter estimates from these ten fits. As the analysis of the individual and population distributions for $\log(\beta)$ and $\delta$ is unchanged from Figure 5, we do not include it here. Rather, we focus on the identifiability of the correlated parameters $\log_{10}T_{0}$ and $\log_{10}p$. In Figure 6A, we show the individual estimates for these parameters for each of the 10 best fits. Then, for each of the 15 individuals, we plot the individual estimates for $\log_{10}T_{0}$ against $\log_{10}p$ and note the clear positive correlation for the 15 individuals in each fit, corresponding to a postive correlation between the individual random effects, but an overall negative correlation when considering each of the 10 fits. We further note that the estimates for each fit do not significantly overlap in $\log_{10}T_{0}\times\log_{10}p$ space. Accordingly, the MMD two-dimensional test identifies significant differences between nearly all pairs of two-dimensional estimates of $(\log_{10}T_{0},\log_{10}p)$, with the blue squares in Figure 6A representing pairs of estimates with $p<0.05$. This analysis suggests that these two parameters may not be identifiable at the population level, even when allowing for a correlation between them.
We next consider the population distributions. Once again, we only show the distributions for $\log_{10}T_{0}$ and $\log_{10}p$ in Figure 6B. Here, we observe the univariate parameter distribtuions reflect the negative correlation structure, with high estimates for $\log_{10}p$ corresponding to lower estimates of $\log_{10}T_{0},$ and vice versa. The lack of overlap between the bivariate parameter distributions is reflected in the relatively low overlapping index, and thus large total variation distance, calculated between the population distribution for each pair of fits. Consequently, our analysis at the individual and population level indicates that these two parameters are unidentifiable.
Figure 6: Practical identifiability of the standard viral dynamics model with a correlation between $\log p$ and $\log T_{0}$ Panel A shows the distribution of the 15 individual estimates for the model parameters $\log_{10}(\beta),\log_{10}(T_{0}),\log_{10}(p)$ and $\delta$ for each of the 10 best population fits as dots within each violin plot. Panel B shows the $p$-value corresponding to the pairwise Kolmogorov-Smirnov test between all 10 best fits for each of the four model parameters. Here, $p$-values below the significance threshold of $0.05$ are plotted as blue squares. Panel C shows the 10 estimated population distributions for each of the four model parameters. Panel D shows the overlapping index between each of the pairwise comparisons of the 10 best fits. Here, for each of the $k=1,2,3,4$ model parameters, pairs of fits $(i,j)$ with an overlapping index $o_{i,j,k}>0.5$ are represented by the shades of violet, while pairs with $o_{i,j,k}<0.5$ are represented by shades of orange. The 10 best fits are distinguished by their colors for the distributions shown in Panels A and C.
@@ 4 Discussion
Nonlinear mixed-effects models are extensively used within the pharmacokinetics modelling community [85, 44, 15] and hierarchical models are increasingly common throughout mathematical biology [41, 29, 8]. Although this modelling framework benefits from simultaneously considering data from multiple individual participants when estimating parameters, existing approaches to study the identifiability of model parameters relies on a parametric interpretation of the population parameter distributions. Here, we have presented an alternative approach to studying the practical identifiability of a nonlinear mixed-effects model through a nonparametric approach. As we do not consider the parametric representation of the population parameter distributions, our approach is immediately applicable in the case where modellers consider different parameter distributions, or in the fully Bayesian case. Specifically, rather than studying when the parametric representation of two distributions are precisely equal, we focus on determining when two distributions are statistically different.
@@
We illustrated our approach by analysing the identifiability of two important models in pharmacology and viral dynamics. We showed that the Friberg model of chemotherapy-induced neutropenia is identifiable within a hierarchical framework without imposing that $k_{prol}=k_{tr}.$ Thus, our work demonstrates that considering multiple individuals allows for the Friberg model to be parameterized without assuming that $k_{tr}=k_{prol}$ [25, 5]. Conversely, our results indicate that the standard viral dynamics model remains unidentifiable, even when considering multiple individuals. This population level analysis is consistent with earlier work that considered the identifiability of the standard viral dynamics model when fitting viral load data from a single individual [87, 7]. Consequently, the standard viral dynamics model is an example of where considering multiple individuals does not permit the identification of additional model parameters. Further, our analysis of the standard viral dynamics model indicates that a multi-start fitting approach, as suggested by Duchesne et al., [23], may not be sufficient to determine a unique global minimum, as we found over 10 local minima of the log-likelihood that have nearly identical goodness-of-fits. These examples are intended to illustrate the proposed nonparametric framework for assessing practical identifiability. Consequently, we focus on evaluating identifiability given the available data, rather than investigating how trial design factors, such as sample size or sampling frequency, might influence identifiability, although these are important questions for future work.
We have proposed a nonparametric approach to studying practical identifiability within the framework of nonlinear mixed-effects models. While our proposed approach is simple, it has some limitations. For example, our proposed method is computationally demanding as we utilize a multi-start approach to seek the global optima of the log-likelihood. Further complicating this multi-start approach, the parameter estimation problem may be high-dimensional and the objective function is not necessarily convex. Consequently, it is not a priori apparent how many runs are necessary to find the global optimum for a given model. Furthermore, we focused on characterizing the population distribution of individual parameters, rather than multivariate joint distributions that would arise from an underlying covariate structure, although we illustrated how our approach extends to the multivariate case in the second example. Nevertheless, there are several areas for future work from both theoretical and practical perspectives. For example, we focused on the identifiability of model parameters for a fixed error model. However, it is possible that using a different error model, which could result in the existence of different local optima, may influence the conclusions of this identifiability analysis [65, 48, 58]. Further, this proposed approach to practical identifiability could be implemented within the convergence assessment framework of existing software packages [78, 8] or the profile likelihood approach to parameter identifiability [70].